-8 is less than or equal to 3r-1 is less than or equal to 1
Found 2 solutions by stanbon, okpuente:
Answer by stanbon(75887)

-8 is less than or equal to 3r-1 is less than or equal to 1
-------
-8 <= 3r-1 <= 1
---
Add 1 along the line to get:
-7 <= 3r <= 2
------
Divide by 3 along the line:
-7/3 <= r <= 2/3
======================
Cheers,
Stan H.
==========

Answer by okpuente(1)

-8<= 3r -1<=1
+1 +1 +1
-7x <=3r <=2
-7x/3 3r/3 2/3
-7/3<= r <=2/3